# Orthogonal Block Kaczmarz Algorithm Based on Preprocessing Technology

### v3

@@ 5.1 Numerical simulation
$RSE=\frac{\left\|x_{j}-x_{*}\right\|_{2}}{\left\|x_{*}\right\|_{2}}.$
@@
We also provided the basic information of the test matrix, including matrix dimension $\left(m\times m\right)$, matrix condition number, and matrix density, as detailed in Table 2. The definition of matrix density is
@@
$Name$ $Kind$ $m$ $Non~{}Num$ $Cond$ $density\left(\%\right)$
$poli$ $Economic~{}Problem$ 4 00840084\,0084 008 8 18881888\,1888 188 3.115e+02 0.051
$muu$ $Structural~{}Problem$ 7 01270127\,0127 012 170 134170134170\,134170 134 7.654e+01 0.337
$poli3$ $Economic~{}Problem$ 16 9551695516\,95516 955 37 8493784937\,84937 849 5.253e+02 0.013
$ex29$ $Computational~{}Fluid~{}Dynamics~{}Problem$ 2 87028702\,8702 870 23 7542375423\,75423 754 7.515e+02 0.288
$kim1$ $2D/3D~{}Problem$ 38 4153841538\,41538 415 933 195933195933\,195933 195 9.863e+03 0.063
$bodyy4$ $Structural~{}Problem$ 17 5461754617\,54617 546 121 550121550121\,550121 550 8.059e+02 0.039
$qpband$ $Optimization~{}Problem$ 20 0002000020\,00020 000 45 0004500045\,00045 000 6.436e+00 0.011
$rajat07$ $Circuit~{}Simulation~{}Problem$ 14 8421484214\,84214 842 63 9136391363\,91363 913 7.894e+02 0.029
$tols4000$ $ComputationalFluidDynamics~{}Problem$ 4 00040004\,0004 000 8 78487848\,7848 784 2.356e+07 0.055
$obstalae$ $Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
$spmsrtls$ $Statistical/Mathematical~{}Problem$ 29 9952999529\,99529 995 229 947229947229\,947229 947 1.416e+05 0.026
$blckhole$ $Structural~{}Problem$ 2 13221322\,1322 132 14 8721487214\,87214 872 4.167e+03 0.327
$linverse$ $Statistical/Mathematical~{}Problem$ 11 9991199911\,99911 999 95 9779597795\,97795 977 3.947e+03 0.067
$torsion1$ $Duplicate~{}Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
$polilarge$ $Economic~{}Problem$ 15 5751557515\,57515 575 33 0743307433\,07433 074 2.983e+01 0.014
$jagmesh4$ $2D/3D~{}Problem$ 1 44014401\,4401 440 9 50495049\,5049 504 1.516e+04 0.458
$bcsstm39$ $Structural~{}Problem$ 46 7724677246\,77246 772 46 7724677246\,77246 772 8.271e+03 0.002
$crystm03$ $Materials~{}Problem$ 24 6962469624\,69624 696 583 770583770583\,770583 770 2.640e+02 0.096
$chem97ztz$ $Statistical/Mathematical~{}Problem$ 2 54125412\,5412 541 7 36173617\,3617 361 2.472e+02 0.114
$wathen100$ $Random~{}2D/3D~{}Problem$ 30 4013040130\,40130 401 471 601471601471\,601471 601 5.816e+03 0.051
@@
Since the GRBK, GREBK(k) and RBK(k) algorithms involve only one block per iteration, yet the aRBK and POBK algorithms involve all the blocks of the matrix per iteration, a direct comparison of the number of iterative steps (IT) is not fair. As can be seen from Table 3, even with $k$ times the number of iteration steps of the POBK algorithm($k<20$), it is still far superior to the GRBK, GREBK(k) and RBK(k) algorithms. As for the computation time, the POBK algorithm is much faster than the randomized chunking algorithms Kaczmarz (GRBK, aRBK) and the k-means direct chunking Kaczmarz (GREBK(k), RBK(k)) algorithms, and we give a comparison of the average running time of the different algorithms for some of the matrices in Figures 5, 6 and 7. It is worth noting that for the matrices over complex field, such as kim1 matrices, the GRBK, GREBK(k), and RBK(k) algorithms encountered problems during the chunking process and could not continue to run. However, the POBK algorithm can run smoothly and has an ideal convergence rate, indicating that it is not only suitable for the real field but also for the complex field.
From the convergence analysis in Section 4, it can be seen that the convergence speed of the POBK algorithm is related to the maximum and minimum non-zero singular values of the blocks and the angle of the hyperplane. We analyzed the matrix structure, and found that other matrices either have smaller ”bandwidth” after preprocessing to accelerate the convergence of the algorithm, or the gap between the largest and smallest non-zero singular values after preprocessing the chunks becomes smaller to accelerate the convergence of the algorithm. However, wathen100 is a very special matrix. Its original matrix is presented as a banded form, and the ”bandwidth” of the RCM algorithm after preprocessing is slightly larger than the original matrix bandwidth (see Figure 8). At the same time, the difference between the maximum and the minimum non-zero singular values after the chunking is also larger (the minimum is 6, and the maximum can reach 40), which will undoubtedly also make the POBK algorithm converge slowly. As a result, the wathen100 matrix has a smaller number of iterations, but the iteration time is comparable to that of the GRBK algorithm. In addition, we find that there is almost no difference in the running time of the wathen100 matrix between preprocessing with the RCM algorithm and without reordering, which reinforces our idea. In this case (sparse matrices with a narrow banding structure), using the original matrix directly may be a better choice. This is also in line with our theme that the POBK algorithm presents good convergence for sparse matrices with scattered arrangements.
Table 3:   IT and CPU performance of 20 test matrices under different algorithms.
$Name$ $GRBK$ $RBK(k)$ $GREBK(k)$ $aRBK$ $POBK$
$poli$ IT 74 5667456674\,56674 566 38 6363863638\,63638 636 51 2595125951\,25951 259 1 72117211\,7211 721 $28$
@@
$muu$ IT $606$ $176$ $269$ $13$ $1$
CPU $2.91$ $7.84$ $2.44$ $3.51$ $0.22$
$poli3$ IT 74 1577415774\,15774 157 57 8185781857\,81857 818 67 8766787667\,87667 876 $899$ 1 03410341\,0341 034
CPU $136.48$ $338.26$ $125.88$ $121.63$ $7.38$
$ex29$ IT $253$ $10$ $116$ $7$ $8$
CPU $0.19$ $1.37$ $0.20$ $0.51$ $0.10$
$kim1$ IT $NAN$ $NAN$ $NAN$ $124$ $52$
CPU $NAN$ $NAN$ $NAN$ $299.93$ $18.27$
$bodyy4$ IT $Inf$ $Inf$ 116 951116951116\,951116 951 18 3251832518\,32518 325 $112$
CPU $NAN$ $NAN$ $807.00$ $3574.54$ $12.97$
$qpband$ IT $627$ $335$ $292$ $11$ $1$
CPU $1.28$ $194.37$ $4.33$ $3.36$ $1.01$
$rajat07$ IT $Inf$ 242 495242495242\,495242 495 234 281234281234\,281234 281 6 89668966\,8966 896 1 08010801\,0801 080
CPU $NAN$ $3094.14$ $3074.96$ $944.40$ $57.52$
$tols4000$ IT $491$ $10$ $927$ $11$ $1$
CPU $0.29$ $4.63$ $0.49$ $0.25$ $0.04$
$obstalae$ IT 37 4063740637\,40637 406 8 46784678\,4678 467 5 02150215\,0215 021 $339$ $127$
CPU $158.43$ $1108.05$ $62.96$ $195.34$ $18.72$
$spmsrtls$ IT $Inf$ $Inf$ $Inf$ $Inf$ $2$
CPU $NAN$ $NAN$ $NAN$ $NAN$ $2.24$
$blckhole$ IT 366 738366738366\,738366 738 22 2132221322\,21322 213 206 208206208206\,208206 208 2 30823082\,3082 308 1 05510551\,0551 055
CPU $456.49$ $32.55$ $261.30$ $151.08$ $10.10$
@@
$torsion1$ IT 37 4243742437\,42437 424 1 19411941\,1941 194 5 04350435\,0435 043 $386$ $132$
CPU $166.10$ $966.06$ $147.52$ $209.02$ $18.57$
$polilarge$ IT $Inf$ 388 865388865388\,865388 865 363 $63 7$ 69776977\,6977 697 5 12851285\,1285 128
CPU $NAN$ $1002.59$ $813.27$ $975.98$ $20.40$
$jagmesh4$ IT $Inf$ 254 433254433254\,433254 433 410 263410263410\,263410 263 7 10471047\,1047 104 2 87328732\,8732 873
CPU $NAN$ $272.82$ $533.79$ $292.11$ $10.40$
$bcsstm39$ IT $20$ $20$ $20$ $4$ $1$
CPU $0.10$ $5189.52$ $19.14$ $4.09$ $0.04$
$crystm03$ IT 4 50345034\,5034 503 $728$ 1 29812981\,2981 298 $29$ $1$
CPU $41.89$ $645.50$ $24.83$ $29.81$ $2.15$
$chem97ztz$ IT 8 72787278\,7278 727 5 34653465\,3465 346 3 57435743\,5743 574 $20$ $9$
CPU $3.88$ $5.08$ $2.03$ $0.61$ $0.04$
$wathen100$ IT 3 16231623\,1623 162 3 52535253\,5253 525 4 96049604\,9604 960 $49$ $150$
C PU $20.78$ $2193.32$ $36.76$ $30.46$ $20.57$
@@
Figure 5:  Average computational time for test matrices $poli3$ and $polilarge$ to reach convergence for different iteration methods.
Figure 6:  Average computational time for test matrices $jagmesh4$ and $rajat07$ to reach convergence for different iteration methods.
Figure 7:  Average computational time for test matrices $muu$ and $chem97ztz$ to reach convergence for different iteration methods.
Figure 8:   Comparison of matrix wathen100 before and after RCM preprocessing.
5.2 The Number of matrix blocks $k$
@@
When the matrix dimension is large, partitioning the matrix is one of the effective ways to improve the convergence of the Kaczmarz algorithm. According to Equation (4.14), the convergence speed of the POBK algorithm is directly related to the cosine of the angle between row blocks. When the number of dimensions $m$ increases, $m>2$. Here, the pinch angle $\theta$ involves both mutually orthogonal blocks and non-mutually orthogonal blocks. When we place mutually orthogonal blocks in a category, the cosine of the angle between these mutually orthogonal blocks will be very small or even close to 0. At this point, it is important to pay attention to the non-mutually orthogonal blocks, i.e., the number of nonzero elements in the cosine table, and the size of these nonzero elements for the rate of convergence of the POBK algorithm to be critical, which also determines the number of matrix chunks. Here we define two variables: the Orthogonality Proportion $zn$ and the Non-orthogonality Severity $nn$ to serve as selection criteria for the number of chunks.
@@
We take the ex29 and wathen100 matrices as examples. For the ex29 matrix, from Figure 9, $zn$ increases as $k$ increases, and $nn$ is small at $k=2,3,5$, but the value of $zn$ at $k=2,3$ is too small, so it is appropriate to choose $k=5$, and the lowest point of the running time box-and-line diagram is also at $k=5$ from Figure 11. For the wathen100 matrix, from Figure 10, $zn$ obtains its maximum value at $k=2$, and it happens that $nn$ also obtains its minimum value at $k=2$, so the choice of $k=2$ is appropriate, and the lowest point of the running time box-and-line plot is also at $k=2$ from Figure 11.
Figure 9: $zn$ and $nn$ values of the ex29 matrix for different values of $k$.
@@
Figure 11:  Boxplots of runtime for ex29 and muu matrices for different values of $k$.
5.3 Stability of the POBK algorithm
The k-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and efficiency, but it requires a given parameter $k$. The difference in the results of each clustering of the k-means algorithm is reflected in numerical experiments as the block Kaczmarz algorithm for k-means clustering, where there is a significant difference in the iteration time even with the same value of $k$, which can lead to the algorithm’s stability is not very good. In our proposed method, given a test matrix $A$, which is preprocessed by the RCM algorithm as an approximate banded matrix $\widetilde{A}$, and $\widetilde{A}$ is unique, and since we perform sequential uniform chunking, the form of the chunks is fixed once $k$ is determined, and thus the algorithm is stable for the same $k$ value. Numerical experiments also show that the algorithm is stable and there is no significant difference in the iteration time for the same value of $k$. Taking the ex29 matrix as an example, Figure 12 shows the iteration time changes of GRBK, RBK(k), GREBK(k), aRBK and POBK algorithms after 10 iterations with the same $k$. It can be clearly seen that the POBK algorithm not only outperforms the other three algorithms in terms of computation time, but also performs better in terms of stability. The stronger stability of the POBK algorithm can also be seen from the POBK runtime boxplot in Section 5.3.

### v4

@@ 3.5 Numerical experiments
$RSE=\frac{\left\|x_{j}-x_{*}\right\|_{2}^{2}}{\left\|x_{*}\right\|_{2}^{2}}.$
@@
We also provided the basic information of the test matrix, including matrix dimension $\left(m\times m\right)$, matrix condition number, and matrix density, as detailed in Table 3. The definition of matrix density is:
@@
$Name$ $Kind$ $m$ $Nonzeros$ $Cond$ $density\left(\%\right)$
$poli$ $Economic~{}Problem$ 4008 8188 3.115e+02 0.051
$muu$ $Structural~{}Problem$ 7012 170134 7.654e+01 0.337
$poli3$ $Economic~{}Problem$ 16955 37849 5.253e+02 0.013
$ex29$ $Computational~{}Fluid~{}Dynamics~{}Problem$ 2870 23754 7.515e+02 0.288
$kim1$ $2D/3D~{}Problem$ 38415 933195 9.863e+03 0.063
$qpband$ $Optimization~{}Problem$ 20000 45000 6.436e+00 0.011
$rajat07$ $Circuit~{}Simulation~{}Problem$ 14842 63913 7.894e+02 0.029
$blckhole$ $Structural~{}Problem$ 2132 14872 4.167e+03 0.327
$linverse$ $Statistical/Mathematical~{}Problem$ 11999 95977 3.947e+03 0.067
$torsion1$ $Duplicate~{}Optimization~{}Problem$ 40000 197608 4.099e+01 0.012
$polilarge$ $Economic~{}Problem$ 15575 33074 2.983e+01 0.014
$jagmesh4$ $2D/3D~{}Problem$ 1440 9504 1.516e+04 0.458
$bcsstm39$ $Structural~{}Problem$ 46772 46772 8.271e+03 0.002
$crystm03$ $Materials~{}Problem$ 24696 583770 2.640e+02 0.096
$chem97ztz$ $Statistical/Mathematical~{}Problem$ 2541 7361 2.472e+02 0.114
@@
Table 4 gives the number of iterations and computation time for these test matrices under different algorithms. Since the RBK, GREBK(k) and RBK(k) algorithms involve only one block of data per iteration, whereas the aRBK and OBK-RCM algorithms involve all the blocks of data of the matrix per iteration, it is not fair to directly compare the number of iteration steps (IT) of the different algorithms. As can be seen from table 4, even if the OBK-RCM algorithm multiplies the number of iteration steps by $k$ times ($k<20$), it still outperforms the RBK, GREBK(K), and RBK(k) algorithms. In terms of computation time, the OBK-RCM algorithm is much faster than the randomized block algorithm Kaczmarz (RBK, aRBK) and the K-means direct block algorithm Kaczmarz(GREBK(k), RBK(k)). It is worth noting that for the matrix kim1 in the complex domain, the RBK, GREBK(k), and RBK(k) algorithms encountered problems during the chunking process and therefore could not continue to run. However, the OBK-RCM algorithm runs smoothly and achieves the desired convergence rate. This shows that it is not only applicable to the real domain but also to the complex domain.
Table 4:   IT and CPU performance of 15 test matrices under different algorithms
$Name$ $RBK$ $RBK(k)$ $GREBK(k)$ $aRBK$ $OBK-RCM$
$poli$ IT 74566 38636 51259 1721 28
@@
$muu$ IT 160 53 269 13 1
CPU 8.44 5.38 2.44 3.51 0.22
$poli3$ IT 26621 6615 18902 1140 1034
CPU 104.62 125.98 48.07 103.04 7.38
$ex29$ IT 123 10 116 7 8
CPU 0.65 1.37 0.20 0.51 0.10
$kim1$ IT NAN NAN NAN 118 52
CPU NAN NAN NAN 276.98 18.27
$qpband$ IT 1687 43 292 11 1
CPU 3.31 12.33 4.33 3.36 1.01
$rajat07$ IT Inf 38150 42957 3478 1080
CPU NAN 2266.82 1889.06 585.08 57.52
$blckhole$ IT 48327 22213 198250 2308 1055
CPU 194.91 32.55 188.74 151.08 10.10
@@
$torsion1$ IT 37424 533 845 386 132
CPU 166.10 119.36 32.65 209.02 18.57
$polilarge$ IT 125195 109525 70322 11102 5128
CPU 411.08 257.46 149.94 625.90 20.40
$jagmesh4$ IT 385621 125814 73429 7104 2873
CPU 1173.66 130.63 96.18 292.11 10.40
$bcsstm39$ IT 50 2 10 4 1
CPU 0.16 122.52 9.30 4.09 0.04
$crystm03$ IT 4503 21 196 29 1
CPU 41.89 36.45 16.20 29.81 2.15
$chem97ztz$ IT 8727 2037 3574 20 9
CPU 3.88 3.64 2.03 0.61 0.04
From the Table 4 and Figure 5, it can be seen that the OBK-RCM algorithm is able to complete the computation within a smaller number of iterations and a shorter CPU time, which exhibits extremely high computational efficiency and stability. This indicates that the OBK-RCM algorithm has significant advantages in dealing with large-scale linear systems with sparse square matrices, especially when the structure of the sparse matrices is more decentralized.
Theoretically, the OBK-RCM algorithm has a better convergence rate by placing $x=P^{T}\widetilde{x}$ inside the loop, which implies that the OBK-RCM algorithm still has potential.
@@
Figure 5:  Average computational time for test matrices $poli$ and $torsion1$ to reach convergence for different iteration methods.
Experimental results (Table 4, Figures 5) demonstrate OBK-RCM’s dual superiority:
  • 1.
    Accelerated Convergence: Requires 50–90% fewer iterations than RBK/GREBK(k) across matrices with dispersed nonzero patterns (e.g., blckhole: 1,055 vs 198,250 iterations).
  • 2.
    Runtime Efficiency: Achieves 10–50 times CPU time (up to several hundred) reduction by eliminating costly K-means clustering and residual sampling (Section 3.3).
This performance stems from RCM’s banded structure optimization and Oclass/Nclass dynamic partitioning—strategies that jointly maximize hyperplane orthogonality while minimizing small-angle configurations, as quantified by Equation (3.16).
3.6 Select the optimal number of blocks of the matrix
As mentioned earlier, for sparse linear systems, blocking is one of the effective means to improve the convergence of kaczmarz algorithm. Taking the $crystm03$ matrix as an example, Figure 7 shows IT and CPU required for the matrix to reach convergence for different values of $k$. It can be seen that when $k=3$, the OBK-RCM algorithm has the optimal number of iterations and running time.
Figure 6:  IT and CPU required for matrix $crystm03$ to reach convergence for different values of $k$.
If $k$ is too large, the dimension of each block becomes smaller, and although the projection computation cost is reduced, the orthogonality between rows of blocks is reduced (i.e., the angle between blocks increases), which slows down convergence; if $k$ is too small, the dimension of a single block is too large, the computational complexity within the block rises and the accelerating effect of the hyperplane orthogonality on the speed of convergence cannot be fully utilized. A proper $k$ maximizes the minimum singular value and thus increases the convergence speed.
@@ 3.6 Select the optimal number of blocks of the matrix
According to Equation (3.16), the convergence speed of the OBK-RCM algorithm is directly related to the angle between rows of blocks, where the angle $\theta$ involves both mutually orthogonal and non-mutually orthogonal blocks. When mutually orthogonal blocks are grouped in a category, the cosine of the angle between these mutually orthogonal blocks will be very small or even close to 0. At this point, it is important to focus on the non-mutually orthogonal blocks, i.e., the number of nonzero elements in the cosine table, and the size of these nonzero elements for the rate of convergence of the OBK-RCM algorithm to be crucial to the number of matrix blocks. Here, two variables are defined: the orthogonality ratio $zn$ and the degree of non-orthogonality $nn$ to be used as selection criteria for the number $k$ of chunks.
@@
Theoretically, the optimal $k$ should simultaneously maximize $zn$ and minimize $nn$.
@@
We take the $muu$ matrices as examples. For the $muu$ matrix, from Figure 7, $zn$ obtains its maximum value at $k=2$, and it happens that $nn$ also obtains its minimum value at $k=2$, so the choice of $k=2$ is appropriate, and the lowest point of the running time box-and-line plot is also at $k=2$ from Figure 8.
@@
Figure 8:  Boxplots of runtime for $muu$ matrices for different values of $k$.
3.7 Determination of threshold $thr$
According to the previous definition, the threshold $thr$ is a key parameter in determining the orthogonality of matrix rows and blocks. Too large or too small a threshold value will reduce the convergence speed of the OBK-RCM algorithm. Therefore, a reasonable choice of the threshold $thr$ requires a trade-off between convergence speed and computational efficiency.
In this experiment, in order to balance the computational time and computational effectiveness, the running times of typical matrices ($ex29$, $blckhole$, $jagmesh4$ and $linverse$) with different $thr$ values are compared, and the $thr$ value with optimal performance for most of the matrices is selected as the global threshold setting. Setting $thr1$, $thr2$, $thr3$, $thr4$, $thr5$, and $thr6$ as $0.5$, $0.2$, $0.1$, $0.05$, $0.02$, and $0.01$, respectively, it can be seen from Table 5 that setting $thr5$ (i.e., setting $thr=0.02$) as the global threshold is worthwhile.
Table 5:   Comparison of processing matrix bandwidth of different sorting algorithms.
$Name$ $thr1$ $thr2$ $thr3$ $thr4$ $thr5$ $thr6$
$ex29$ 0.1126 0.1097 0.1087 0.1086 0.1012 0.1014
$linverse$ 0.4295 0.4315 0.4281 0.4259 0.4265 0.4281
$blckhole$ 10.4855 10.257 10.1727 10.2027 10.1444 10.2011
$jagmesh4$ 10.5939 10.6184 10.5471 10.5426 6810.5386 10.5355
3.8 Stability of the OBK-RCM algorithm
The K-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and high efficiency, but the results of this algorithm vary each time the clustering is performed, which is reflected in the numerical experiments as the block Kaczmarz algorithm for K-means clustering, where the computation time is very different even for the same $k$ value, which leads to the algorithm’s stability is not very good.